OpenStudy (anonymous):

how to solve this by factoring 2x^2 - 3x = 0

OpenStudy (anonymous):

Take an x out of each term . So we have x(SOMETHING - SOMETHING) = 0

OpenStudy (anonymous):

so is it two separate equations like x(x - 1) = 0 x-2 = 0?

OpenStudy (anonymous):

No it's still one equation.

OpenStudy (anonymous):

oh so x( 2 - 3)? I'm still kind of confused

OpenStudy (anonymous):

If it's x( 2 - 3) , then we have 2x-3x.

OpenStudy (anonymous):

We want \[2x ^{2}\] instead.

OpenStudy (perl):

$$ \Large { 2x^2 - 3x = 0 \\ \iff \\ x (2x-3) = 0 \\ \iff \\ x = 0 ~~or ~~ 2x-3 = 0 } $$

OpenStudy (anonymous):

Oh ok I get it now :) thanks